A loan officer states, "Thousands of dollars can be saved by switching to a 15-year mortgage from a 30-year mortgage." Calculate the difference in payments on a 30-year mortgage loan (APR of 3.5%) vs. a 15-year mortgage (APR of 3.0%). Assume both mortgages are for $100,000 and require monthly payments. Then, calculate the total payments over the entire loan period under each scenario. What is the difference in total loan payments to the lender between the 2 scenarios? (Round the monthly payment amounts to 2 decimal places.)

Group of answer choices

$161,656.09

$40,242.36

$37,351.39

$124,304.70
Answer #1

Calculating Monthly Payment on 30 year loan,

Using TVM Calculation,

PMT = [PV = 100,000, FV = 0, N = 360, I = 0.035/12]

PMT = $449.04

Calculating Monthly Payment on 15 year loan,

Using TVM Calculation,

PMT = [PV = 100,000, FV = 0, N = 180, I = 0.03/12]

PMT = $690.58

Difference in Payment = 360(449.04) - 180(690.58)

Difference in Payment = $37,351.39